A 17-year, semiannual coupon bond sells for $948.63. The bond has a par value of $1000 and a yield to maturity of 7.1 percent. What is the bond's coupon rate? 5.93% 4.94% 6.26% 6.58% 3.29%
Answer #1

Number of periods = 17 * 2 = 34

Rate = 7.1% / 2 = 3.55%

Price of bond = Coupon * [ 1 - 1 / ( 1 + r)n] / r + FV / ( 1 + r)n

948.63 = Coupon * [ 1 - 1 / ( 1 + 0.0355)34] / 0.0355 + 1,000 / ( 1 + 0.0355)34

948.63 = Coupon * 19.565653 + 305.419309

Coupon = 32.874481

Semi annual coupon rate = (32.874481 / 1,000) * 100

Semi annual coupon rate = 3.287448%

Annual coupon rate = 3.287448% * 2

Annual coupon rate = 6.58%